Mol NH_3 are placed in a 10.2 L flask and allowed
Answer #1

            2NH3(g) --------> 3H2(g) + N2(g)

I            3.36                      0             0

C        -2*1.33                 3*1.33     1.33

E         0.7                       3.99          1.33

  

Kc   = [N2][H2]3/[NH3]2

    [NH3] = 0.7/10.2   = 0.068M

   [H2]    = 3.99/10.2   = 0.39M

   [N2] = 1.33/10.2 = 0.13M

Kc   = [N2][H2]3/[NH3]2

      = 0.13*(0.39)3/(0.068)2

     = 0.019773/0.004628 = 4.2724